# FALCONEye: Finding Answers and Localizing Content in ONE-hour-long videos with multi-modal LLMs

### v1

@@ 5.1 FALCON-Bench results
Table 2 presents the performance of baselines, open-source VLMs, and meta-architectures on FALCONBench. As baselines, we include human performance both to establish an upper bound and to analyze human strategies in VAS, which guided our algorithm design (Sec. 8.4). Additionally, we evaluate our two selected LLMs (GPT4o-mini and Qwen2.5-VL) in blind mode (without visual information) to assess their reliance on text-based priors. For open-source VLMs, we test small-scale versions of the main state-of-the-art models, including Apollo and MovieChat, which are specifically designed for long videos. Additionally, we evaluate two well-known meta-architectures, Socratic [50] and VideoAgent [43]. To ensure a fair comparison with our FALCONEye, both meta-architectures are implemented using the same VLM (Qwen2.5-VL) and LLM (GPT4o-mini). The Socratic approach generates captions (768 tokens) from 1-minute clips, while VideoAgent is extended to handle open-ended questions and predict answer localization, resulting in VideoAgent++. Most VLMs that are limited to a small number of sampled frames fail to outperform the LLM-blind baselines in MCQs, suggesting that their advantage over random guessing stems from discarding certain options, rather than actual answer retrieval from visual content. A similar issue arises with Qwen2.5-VL when operating in its default mode (2fps, up to 768 frames), as it downscales frames to $140\times 240$, losing crucial details. Only LLaVA-Video and Apollo show clear improvements over the blind baselines. This trend is not observed in OQs, where models must generate answers independently, validating their reliability. Among meta-architectures, only FALCONEye demonstrates robustness for long-form VAS, achieving 70.0% accuracy in MCQs and 44.7% in OQs with its top-performance configuration, and 64.7% and 41.1%, respectively, in the cost-efficient variant. The latter configuration achieves performance close to the Socratic baseline while maintaining a processing time per question not so far from human response times. We analyze their costs and performances, including exploration baselines, in Table 3. We also include Socratic [Uncaptioned image], which utilizes the short captions (64 tokens) generated by FALCONEye. While baseline exploration algorithms outperform existing meta-architectures, they require significantly more time. Notably, our LLM-guided exploration algorithm reduces this overhead while simultaneously improving performance by efficiently narrowing the search.
5.2 Other benchmarks results
We evaluate FALCONEye on additional VQA benchmarks to compare its performance against state-of-the-art private and public large-scale VLMs. From the benchmarks listed in Tab. 1, we select MLVU [53] and LONGVIDEOBENCH [44], as both were recently published in top-tier conferences, include tasks related to single-detail questions (closely aligned with VAS), and are available within the lmms-eval framework. A summary of these results is presented in Tab. 4.
Although these benchmarks do not fully capture the two key challenges addressed by FALCONEye: one-hour-long videos (these benchmarks instead average only a few minutes per video, where the 256 frames sampled by top VLMs represent a significant portion of the entire video), and open-ended question (OQ) evaluation (these benchmarks primarily assess MCQs)—FALCONEye demonstrates competitive performance against top VLMs, achieving the new state-of-the-art in some MLVU subtasks.
@@ 5.2 Other benchmarks results
Table 5: Some VLMs calibration and performance comparison with GT 1min-length clips in the test split of FALCONBench.
@@
s # Res. Avg. ACE $\downarrow$ CC $\uparrow$ Acc. Sc. ACE $\downarrow$ CC $\uparrow$
@@
Figure 3: Qwen2.5-VL calibration plots with the GT 1min-length clips of the FALCONBench test split. Left: MCQs. Right: OQs.
6 Configuration Analysis
VLMs calibration study.
Since our exploration algorithm relies on answer confidence, here we evaluate if the confidence values are well calibrated. For that, we adopt Reliability Diagrams [12], which group predictions into bins based on confidence and measure the gap (calibration error) between confidence and accuracy for each bin. From these plots, we compute the Average Calibration Error (ACE) and Calibration Count (CC), quantifying the percentage of predictions above a defined confidence threshold, weighted by their accuracy calibration error (1-CE).
Given the poor performance of state-of-the-art VLMs on FALCONBench, we ease the task by providing directly one-minute clips taken from the middle of the ground-truth (GT) temporal interval (Sec. 10.1 of supp. material). For MCQs, confidence computation is straightforward, as only a single token is output (Sec. 3.2). However, for OQs, confidence must be aggregated across the entire sequence of tokens. Similar to calibration studies for LLMs [26], we investigate various aggregation metrics in Sec. 10.2, identifying the geometric average (Eq.1) as the most suitable approach for VLMs. Table 5 presents the calibration analysis for SOTA VLMs, with all methods showing strong calibration in MCQs. For OQs, Qwen2.5-VL significantly outperforms other approaches. The calibration plots for Qwen2.5-VL are shown in Fig. 3. The Section 10.2 of the supplementary material includes deeper calibration analysis.
Table 6: Qwen2.5-VL performance comparison when varying the GT clip length of FALCONBench validation split.
Visual Information MCQs OQs
$\Delta t$ Frames Accuracy Accuracy Score
s # Res. M S T Avg. M S T Avg. M S T Avg.
60 32 $672\times 1204$ 85.0 71.4 86.7 81.0 60.0 60.0 66.7 62.2 3.15 3.54 3.53 3.43
5 2fps $824\times 1462$ 85.0 74.3 80.0 79.8 70.0 62.9 73.3 68.7 3.70 3.71 3.80 3.73
Table 7: FALCONEye ablation study in the OQs of FALCONBench validation split. We vary the VLM [Uncaptioned image] (Qwen2.5-VL [4] and LLaVA-Video [52]), the LLM [Uncaptioned image] (GPT4o-mini, GPT4o and, GPTo3-mini), the zoom-in effect ([Uncaptioned image] or fix resolution), the max. number of evaluated candidate clips [Uncaptioned image] and the confidence threshold [Uncaptioned image]. We report the average accuracy (Acc.) and score (Sc.).
FALCONEye LLM VLM Time OQs
N [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] #Inf. #T #Inf.T s Acc. Sc.
[52] 4o-m 45 0.8 7.0 1.5 41 549.1 17.5 1.13
[Uncaptioned image] [4] 4o-m 45 0.8 3.7 1.7K 21.2 229.2 46.6 2.47
[4] 4o 45 0.8 5.1 1.6K 24.4 274.8 48.8 2.56
[4] o3-m 45 0.8 6.5 3.6K 25.6 315.8 50.4 2.63
[4] 4o-m 45 0.8 4.3 1.7K 23.8 258.1 43.0 2.31
[4] 4o-m 25 0.8 3.4 1.7K 19.7 212.7 45.6 2.45
[Uncaptioned image] [4] 4o-m 10 0.8 2.9 1.8K 17.5 188.3 42.6 2.26
[4] 4o-m 45 0.7 4.0 1.7K 18.6 205.4 45.0 2.44
[4] 4o-m 45 0.9 5.8 1.8K 34.9 375.6 49.5 2.64
Zoom-in effect study.
Qwen2.5-VL can process images of any resolution, dynamically converting them into a variable number of visual tokens. This creates a trade-off between number of frames and frame resolution when processing videos within the context window size limit. This trade-off is analyzed in the validation split of FALCONBench by varying the clip length (Tab. 6). The best configuration per clip-level category is selected for FALCONEye.
@@ 6 Configuration Analysis
Tab. 7 validates some details of our best configurations FALCONEye [Uncaptioned image] and FALCONEye [Uncaptioned image]. While other configurations are more accurate, we must take into the account efficiency. For LLMs choice, it is important to note that this validation split (70Q) experiment costs: GPT4o-mini \$0.1, GPT4o \$2.0, and GPTo3-mini \$3.5.
@@ 7 Conclusions
We present FALCONEye, a novel video agent driven by an efficient exploration algorithm integrating a LLM and a VLM to tackle the challenge of answering open-ended questions in hour-long videos (VAS) with answers located in short temporal windows. We evaluate our meta-architecture using a 7B-size VLM and a cost-efficient LLM (academic resources), and demonstrate that it significantly outperforms sota 7B-size VLMs and similar video agents on FALCON-Bench, a new designed benchmark for the VAS task. Besides, when applied to shorter videos and multiple-choice questions in existing benchmarks, it achieves similar performance on average to top VLMs like GPT-4o.
@@ 8.1 Examples and video source details
  • SoccerNet [16] – 56 videos were selected from this dataset, each averaging 92.4 minutes in duration. These structured soccer match recordings include 389 questions. The mean GT time interval is 61.9 seconds.
  • MovieChat-1K Films [38] – a total of 140 film clips, each 8 minutes long, were selected from the dataset. Clips from the same film were combined to create 24 film segments, with an average duration of 46.4 minutes each. This subset contains 122 questions, with a mean GT time interval of 22.4 seconds.
  • Walking Tours Dataset [40] – 12 high-resolution videos ($3840\times 2160$) were selected from this dataset, averaging 81.3 minutes. These videos capture city tours from an egocentric perspective and include 65 questions. The mean GT time interval is 31.4 seconds.
@@ 8.3 Localization evaluation
FALCONBench requires models to provide an evidence of the answer (short clip) rather than precisely matching the entire clip where the answer is located. To achieve this goal, we leverage The Ground Truth over Union (GToU) metric to compare the predicted and the GT temporal interval. Unlike the commonly used Intersection over Union (IoU) in temporal grounding tasks [13], GToU assigns a score of 1.0 if the predicted interval is entirely within the GT interval, regardless of the degree of overlap. In all other cases, GToU behaves identically to IoU (Figure 6). In mathematical terms,
@@ 8.4 Human experiments
Participants were seated in front a test computer equipped with two monitors, one for watching the video and another for editing a .json file. They were provided with a structured .json file containing the question details in the following format:
Listing 1: JSON question example provided to volunteers.
"id": "WalkingTours-009",
"question": "Whichcarcompanyhasa
servicecenternexttoToyota?",
"video": "Chiang_Mai.mp4",
"options": null,
"type": "Open-endedquestion",
"answer": null,
"temporal_window": null,
"spent_time": null
@@ 9 FALCONEye: more details
In this section we give further implementation details of our FALCONEye video agent explained in Sec. 3.
@@ 9.1 Exploration Algorithm pseudo-code
Algorithm 1 shows the pseudo-code of our FALCONEye exploration algorithm proposed in Sec. 3.2 and which can be visualized in Fig. 2.
@@
Output: Answer $\mathcal{A}$, Confidence $p(\mathcal{A})$, $Clip$ $v\in\mathcal{V}$
@@ 9.2 Prompts
Figure 8: LLM prompts used through FALCONEye algorithm.
@@ 10.1 FALCON-Bench
Given the poor performance of state-of-the-art VLMs on FALCON-Bench, we further evaluated VLMs under a simplified VAS setup. Specifically, we extracted 1-minute clips centered within the ground truth intervals to ensure that the answers were contained within these shorter segments (Table 9). Qwen2.5-VL achieved the best accuracy and score in OQs, and LLaVA-Video in MCQs. However, even after reducing the search space from an hour-long video to just one minute, the performance remains relatively low. To address this, we took an additional step and tested the VLMs by providing a single frame that contains the answer, which is always within the ground truth interval defined by FALCONBench (Table 10). These results represent the maximum performance our FALCONEye agent could achieve with each VLM, assuming the ground truth frame is the optimal frame for answering the question. In this evaluation, GPT-4o achieved the highest performance, followed by Qwen2.5-VL.
Table 9: Model performance on FALCONBench test split with GT 1min-length clips containing the answer. We report average accuracy for MCQs and OQs across MovieChat (M), SoccerNet (S), WalkingTours (T), and overall (Avg.).
@@
Open-Source VLMs with Multi-Image Support
@@
mPLUG-Owl3 [47] 57.6 39.2 21.1 26.0 28.7 2.24 1.20 1.58 1.67
LLaVA-OV [17] 72.2 45.0 28.5 46.0 39.8 2.46 1.57 2.42 2.15
@@
Open-Source VLMs designed for Videos
VideoChat2-HD [21] 27.2 18.6 12.1 22.0 17.5 1.24 0.67 1.34 1.08
Video-LLAVA [22] 34.0 19.6 6.49 18.0 14.6 1.20 0.46 1.14 0.93
LLaVA-Video [19] 79.2 61.7 44.0 52.0 52.6 3.31 2.38 2.74 2.81
Open-Source VLMs specific for long videos
MovieChat-OV [38] 44.5 5.88 18.3 14.0 12.7 0.49 1.02 0.82 0.77
Apollo [54] 71.3 50.9 31.3 48.0 43.4 2.77 1.79 2.62 2.39
@@
mPLUG-Owl3 [47] 82.9 59.8 53.1 58.0 65.2 3.25 2.89 3.18 3.10
LLaVA-OV [17] 84.4 40.1 10.4 54.0 34.8 2.02 0.53 2.80 1.78
@@
LLaMa3.2M [1] 85.9 69.6 64.9 70.0 68.1 3.68 3.27 3.70 3.55

### v2

@@ 5.1 FALCON-Bench results
Table 4 presents the performance of baselines, open-source VLMs, and meta-architectures on FALCON-Bench. As baselines, we include human performance both to establish an upper bound and to analyze human strategies in VAS, which guided our algorithm design (Sec. 8.4). Additionally, we evaluate our two selected LLMs (GPT4o-mini and Qwen2.5-VL) in blind mode (without visual information) to assess their reliance on text-based priors. For open-source VLMs, we test small-scale versions of the main state-of-the-art models, including Apollo and MovieChat, which are specifically designed for long videos. We also evaluate three video agents, Socratic [zeng2022socratic], LifelongMemory [wang2024lifelongmemory] for VQA and VideoAgent [wang2024videoagent]. To ensure a fair comparison with our FALCONEye, these meta-architectures are implemented using the same VLM (Qwen2.5-VL) and LLM (GPT4o-mini). The Socratic approach generates captions (768 tokens) from 1-minute clips, while VideoAgent is extended to handle open-ended questions and predict answer localization, resulting in VideoAgent [wang2024videoagent]++OQ. Most VLMs that are limited to a small number of sampled frames fail to outperform the LLM-blind baselines in MCQs, suggesting that their advantage over random guessing stems from discarding certain options, rather than actual answer retrieval from visual content. A similar issue arises with Qwen2.5-VL when operating in its default mode (2fps, up to 768 frames), as it downscales frames to $140\times 240$, losing crucial details. Only LLaVA-Video, Apollo, and ReKV show clear improvements over the blind baselines. This trend is not observed in OQs, where models must generate answers independently, validating their reliability. Among meta-architectures, only FALCONEye demonstrates robustness for long-form VAS, achieving 70.0% accuracy in MCQs and 44.7% in OQs with its top-performance configuration, and 64.7% and 41.1%, respectively, in the cost-efficient variant. The latter configuration achieves performance close to the Socratic baseline while maintaining a processing time per question not so far from human response times. We analyze their costs and performances, including exploration baselines, in Table 3. We also include Socratic-Short, which utilizes the short captions (64 tokens) generated by FALCONEye. While baseline exploration algorithms outperform existing meta-architectures, they require significantly more time. Notably, our LLM-guided exploration reduces this overhead while improving performance by efficiently narrowing the search.
Table 4: Model performance comparison on a VQA benchmarks with medium-duration videos — MLVU (720s), covering a diverse set of tasks beyond Single-Detail question types. FALCONEye surpasses comparable size/cost agents and GPT-4o at SD tasks.
MLVU Multiple-Choice Generation
Single-Detail Non-Single-Detail M-Avg SD NSD G-Avg
Model Name NQA ER PQA Avg. TR AR AO AC Avg. SSC VS
[…]
GPT-4o [openai2023gpt4v] 42.9 47.8 57.1 49.2 83.7 68.8 46.2 35.0 58.4 54.5 6.80 4.94 5.87
[…]
VideoAgent [wang2024videoagent] 28.3 30.1 28.0 28.8 72.5 23.0 21.4 11.6 32.1 30.7 4.61 3.35 3.98
FALCONEye-Pro 51.6 45.2 52.0 49.6 80.2 30.7 27.1 3.33 35.3 41.4 5.04 3.40 4.22
5.2 FALCONEye generalization evaluation
To explore FALCONEye generalization beyond our benchmark, we evaluate it out-of-the-box on a standard VQA setting: shorter videos and a broader mix of question types beyond single-detail (SD). With this goal, we select MLVU [zhou2024mlvu], the most recent publicly available benchmark covering various VQA subtasks, not only SD questions. A summary of these results is presented in Table 5.1.
While MLVU does not reflect the two core challenges tackled by FALCONEye—namely, one-hour-long videos (here, video durations are only a few minutes) and open-ended questions (MLVU primarily includes multiple-choice questions)—it still offers valuable insights. Notably, shorter videos benefit standard VLMs like GPT-4o, as their frame sampling strategies (e.g., 256 frames) cover a large portion of the content, reducing the need for exploration.
[…]
6 Configuration Analysis
VLMs confidence calibration study.
Since our exploration algorithm relies on answer confidence, here we evaluate if the confidence values are well calibrated, i.e., the confidence correlates with the accuracy. For that, we adopt Reliability Diagrams [degroot1983comparison], which group predictions into bins based on confidence and measure the gap (calibration error) between confidence and accuracy for each bin. From these plots, we compute the Average Calibration Error (ACE) and Calibration Count (CC), quantifying the percentage of predictions above a defined confidence threshold, weighted by their accuracy calibration error (1-CE).
Given the poor performance of state-of-the-art VLMs on FALCON-Bench, we ease the task by providing directly one-minute clips taken from the middle of the ground-truth (GT) temporal interval (Sec. 10.1 of supp. material). For MCQs, confidence computation is straightforward, as only a single token is output (Sec. 3.2). However, for OQs, confidence must be aggregated across the entire sequence of tokens. Similar to calibration studies for LLMs [liu2023litcab], we investigate various aggregation metrics in Sec. 10.3, identifying the geometric average (Eq.1) as the most suitable approach for VLMs. Table 5 presents the calibration analysis for SOTA VLMs, with all methods showing strong calibration in MCQs. For OQs, Qwen2.5-VL significantly outperforms other approaches. The Section 10.3 of the supplementary material includes deeper calibration analysis, including the corresponding models’ calibration plots.
@@ 6 Configuration Analysis
Table 5: Calibration and performance analysis for commonly used VLMs on GT 1min-length clips in the test split of FALCON-Bench.
@@
s # Res. Avg. ACE $\downarrow$ $C$ $\uparrow$ Acc. Sc. ACE $\downarrow$ $C$ $\uparrow$
@@
Table 6: FALCONEye study of different meta-architectures. We vary the LLM and VLM used in FALCONEye-Pro (first row) and evaluate in both MCQs and OQs on FALCON-Bench validation split, including number of LLM tokens (#T\#T) used per question.
FALCONEye MCQs OQs
VLM LLM #T\#T Acc. Loc. #T\#T Acc. Sc. Loc.
Qwen2.5-VL-7B GPT-4o-m 8.4K 59.6 27.3 7.6K 46.6 2.47 26.9
LLaVA-Video-7B GPT-4o-m 6.7K 52.2 10.9 10.9K 17.5 1.13 2.71
Qwen2.5-VL-3B GPT-4o-m 11.9K 57.1 25.6 9.4K 22.8 1.24 13.9
Qwen2.5-VL-7B GPT4o 6.5K 59.7 33.4 8.7K 48.8 2.56 23.8
Qwen2.5-VL-7B GPTo3-m 27.8K 63.4 32.3 28.8K 50.4 2.63 29.3
Qwen2.5-VL-7B GPT5 14.0K 69.2 47.2 18.8K 47.1 2.49 35.5
Qwen2.5-VL-7B Gem2.5-Flash Lite 21.0K 56.6 27.0 18.4K 36.1 1.96 21.8
Qwen2.5-VL-7B Gem2.5-Flash 28.4K 69.8 38.2 32.0K 47.3 2.59 36.9
Qwen2.5-VL-7B Gem2.5-Pro 12.3K 69.5 39.6 17.2K 52.3 2.88 41.7
Table 7: Study of FALCONEye confidence estimation method.
FALCONEye MCQs OQs
Confidence Acc. Loc. Acc. Sc. Loc.
No Confidence 42.0 21.7 33.4 1.84 22.0
VLM Verbal Confidence 53.4 24.3 35.4 1.87 23.0
LLM Verbal Confidence 41.2 13.5 34.8 1.82 20.6
Geom.Avg (Ours) 59.6 27.3 46.6 2.47 26.9
@@
We validate FALCONEye’s main features by first illustrating its training-free and model agnostic capabilities in Table 6, showing that current FALCONEye-Pro performance could be easily increased using a more powerful LLM or even VLM. Second, Table 7 shows the significance of using a calibrated confidence estimation of the VLM answers inside our exploration algorithm. To achieve this, we compare our confidence estimation mechanism against those used in related works: verbal confidence estimated together with the answer (by the VLM in our case) [zhi2025videoagent2] and verbal confidence estimated separately from the answer by an LLM given the question, answer and video context [wang2024videoagent]. We also perform a detailed ablation of the different FALCONEye’s exploration algorithm stages to show their contribution, including the zoom-in effect. The details of this study are in Sec. 10.2 of supplementary material.
@@ 7 Conclusions
We present FALCONEye, a novel video agent that integrates a VLM and an LLM via an efficient exploration algorithm, enabling the answering of single-detail, open-ended questions over long-form videos. We evaluate our meta-architecture using a 7B-size VLM and a cost-efficient LLM (academic resources), and demonstrate that it significantly outperforms sota 7B-size VLMs and similar video agents on FALCON-Bench, a new designed benchmark for the Video Answer Search task. Besides, when applied to shorter videos and broader tasks in MLVU benchmark, it surpasses GPT-4o on single-detail tasks while slashing inference cost by roughly an order of magnitude.
@@ 8.1 Examples and video source details
  • SoccerNet [giancola2018soccernet] – 56 videos were selected from this dataset, each averaging 92.4 minutes in duration. These structured soccer match recordings include 389 questions. The mean GT time interval is 61.9 seconds.
  • MovieChat-1K Films [song2024moviechat] – a total of 140 film clips, each 8 minutes long, were selected from the dataset. Clips from the same film were combined to create 24 film segments, with an average duration of 46.4 minutes each. This subset contains 122 questions, with a mean GT time interval of 22.4 seconds.
  • Walking Tours Dataset [venkataramanan2023imagenet] – 12 high-resolution videos ($3840\times 2160$) were selected from this dataset, averaging 81.3 minutes. These videos capture city tours from an egocentric perspective and include 65 questions. The mean GT time interval is 31.4 seconds.
@@ 8.3 Localization evaluation
FALCON-Bench requires models to provide an evidence of the answer (short clip) rather than precisely matching the entire clip where the answer is located. To achieve this goal, we leverage The Ground Truth over Union (GToU) metric to compare the predicted and the GT temporal interval. Unlike the commonly used Intersection over Union (IoU) in temporal grounding tasks [ding2023temporal], GToU assigns a score of 1.0 if the predicted interval is entirely within the GT interval, regardless of the degree of overlap. In all other cases, GToU behaves identically to IoU (Figure 5). In mathematical terms,
@@ 8.4 Human experiments
Participants were seated in front a test computer equipped with two monitors, one for watching the video and another for editing a .json file. They were provided with a structured .json file containing the question details.
@@ Results
Figure 7: Visualization of the Socratic baseline approach together with our three exploration baselines considered to address VAS.
@@ 9 FALCONEye: more details
This section gives further implementation details of our FALCONEye video agent explained in Sec. 3.
@@ 9.1 Exploration Algorithm pseudo-code
Algorithm 1 shows the pseudo-code of FALCONEye exploration algorithm (explained in Sec. 3.2 and Fig. 2).
@@
Output: Answer $\mathcal{A}$, Confidence $p(\mathcal{A})$, ClipClip $v\in\mathcal{V}$
@@ 9.2 Prompts
Figure 8: LLM prompts used in FALCONEye algorithm and and FALCON-Bench.
@@ 10.1 FALCON-Bench
Given the poor performance of state-of-the-art VLMs on FALCON-Bench, we further evaluated VLMs under a simplified VAS setup. Specifically, we extracted 1-minute clips centered within the ground truth intervals to ensure that the answers were contained within these shorter segments (Table 10.1). Qwen2.5-VL achieved the best accuracy and score in OQs, and LLaVA-Video in MCQs. However, even after reducing the search space from an hour-long video to just one minute, the performance remains relatively low. To address this, we took an additional step and tested the VLMs by providing a single frame that contains the answer, which is always within the ground truth interval defined by FALCON-Bench (Table 10.1). These results represent the maximum performance our FALCONEye agent could achieve with each VLM, assuming the ground truth frame is the optimal frame for answering the question. In this evaluation, GPT-4o achieved the highest performance, followed by Qwen2.5-VL.
Table 9: Model performance on FALCON-Bench test split with GT 1min-length clips containing the answer. We report average accuracy for MCQs and OQs across MovieChat (M), SoccerNet (S), WalkingTours (T), and overall (Avg.).
@@
[…]
@@
mPLUG-Owl3 [ye2024mplug] 57.6 39.2 21.1 26.0 28.7 2.24 1.20 1.58 1.67
LLaVA-OV [llava-onevision] 72.2 45.0 28.5 46.0 39.8 2.46 1.57 2.42 2.15
@@
[…]
VideoChat2-HD [li2024mvbench] 27.2 18.6 12.1 22.0 17.5 1.24 0.67 1.34 1.08
Video-LLAVA [lin2023video] 34.0 19.6 6.49 18.0 14.6 1.20 0.46 1.14 0.93
LLaVA-Video [llava-next] 79.2 61.7 44.0 52.0 52.6 3.31 2.38 2.74 2.81
[…]
MovieChat-OV [song2024moviechat] 44.5 5.88 18.3 14.0 12.7 0.49 1.02 0.82 0.77
Apollo [zohar2024apollo] 71.3 50.9 31.3 48.0 43.4 2.77 1.79 2.62 2.39
[…]
FALCONEye-Pro 81.7 63.7 70.0 75.0 66.7 3.39 3.51 3.56 3.49
FALCONEye-Flash 81.9 66.6 69.7 70.0 68.8 3.54 3.62 3.62 3.59
@@
mPLUG-Owl3 [ye2024mplug] 82.9 59.8 53.1 58.0 65.2 3.25 2.89 3.18 3.10
LLaVA-OV [llava-onevision] 84.4 40.1 10.4 54.0 34.8 2.02 0.53 2.80 1.78
@@
LLaMa3.2M [llama3_2] 85.9 69.6 64.9 70.0 68.1 3.68 3.27 3.70 3.55
10.2 Ablation study of exploration algorithm stages
We validate FALCONEye’s exploration algorithm with an ablation study of its different stages, explained in Sec. 3.2, shown in Table 11. It is noted that each stage contributes significantly to achieving the FALCONEye’s superior performance compared to the baselines.
Table 11: FALCONEye ablation study of the exploration algorithm. We compare the time and performance gain that each of the four stages of our exploration algorithm brings (① Pre-processing, ② Reasoning, ③ Evaluation, and ④ Decision, as detailed in Sec. 3). To validate them, we first measure performance when giving the whole video to the VLM and the captions extracted in Stage ① to the LLM. Lately, we validate the stages adding them sequentially and comparing the reasoning stages ② and ④ vs random guess.
FALCONEye MCQs OQs
Exploration Algorithm s Acc. Loc. s Acc. Sc. Loc.
Video $\rightarrow$ VLM 68.9 23.4 0.00 70.1 11.3 0.71 0.00
[…]
+Random+ 174.9 26.1 3.16 165.2 9.92 0.60 4.06
++ 181.2 51.9 22.1 171.6 34.8 1.90 21.2
+++Random 383.2 54.8 26.8 290.8 38.9 2.01 21.9
FALCONEYE:①+②+③+④ 348.7 59.6 27.3 229.2 46.6 2.47 26.9
We also perform a more detailed study of the influence of the zoom-in effect incorporated in our approach. Qwen2.5-VL can process images of any resolution, dynamically converting them into a variable number of visual tokens. This creates a trade-off between number of frames and frame resolution when processing videos within the context window size limit. This trade-off is analyzed in the validation split of FALCON-Bench by varying the clip length (Table 12). The best configuration per clip-level category is selected for FALCONEye.
Table 12: Qwen2.5-VL performance comparison when varying the GT clip length of FALCON-Bench validation split.
Visual Information MCQs OQs
$\Delta t$ Frames Accuracy Accuracy Score
s # Res. M S T Avg. M S T Avg. M S T Avg.
[…]